# SymbolNet: Neural Symbolic Regression with Adaptive Dynamic Pruning

### v2

@@ Abstract
Contrary to genetic programming, the neural network approach to symbolic regression can efficiently handle high-dimensional inputs and leverage gradient methods for faster equation searching. Common ways of constraining expression complexity often involve multistage pruning with fine-tuning, which can result in significant performance loss. In this work, we propose $\tt{SymbolNet}$, a neural network approach to symbolic regression in a novel framework that allows dynamic pruning of model weights, input features, and mathematical operators in a single training process, where both training loss and expression complexity are optimized simultaneously. We introduce a sparsity regularization term for each pruning type, which can adaptively adjust its strength, leading to convergence at a target sparsity ratio. Unlike most existing symbolic regression methods that struggle with datasets containing more than $\mathcal{O}(10)$ inputs, we demonstrate the effectiveness of our model on the LHC jet tagging task (16 inputs), MNIST (784 inputs), and SVHN (3072 inputs). Our approach enables symbolic regression to achieve fast inference with nanosecond-scale latency on FPGAs for high-dimensional datasets in environments with stringent computational resource constraints, such as the high-energy physics experiments at the LHC.
@@ 1 Introduction
Symbolic regression (SR) is a supervised learning method that searches for analytic expressions that best fit the data. Unlike traditional regression methods, such as linear and polynomial regression, SR can model a much broader range of complex datasets because it does not require a pre-defined functional form, which itself is dynamically evolving in the fit.
@@
By expressing models in symbolic forms, SR facilitates human interpretation of the data, enabling the potential inference of underlying principles governing the observed system, in contrast to the opaque nature of black-box deep learning (DL) models. A historical example is Max Planck’s 1900 empirical fitting of a formula to the black-body radiation spectrum [1], known as Planck’s law. This symbolically fitted function not only inspired the physical derivation of the law but also played a key role in the revolutionary development of quantum theory.
Moreover, due to its compact representation compared to most DL models, SR can also be used as a distillation method for model compression. This can accelerate inference time and reduce computational costs, making it particularly valuable in resource-constrained environments.
However, a significant drawback of SR is its inherent complexity. The search space for equations expands exponentially with the number of building blocks (variables, mathematical operators, and constants), making it a challenging combinatorial problem. In fact, finding the optimal candidate has been shown to be NP-hard [2].
@@
Genetic programming (GP) has traditionally been the primary approach to SR. It constructs expressions using a tree representation, where the algebraic relations are reflected in the tree’s structure. The tree’s lowest nodes consist of constants and variables, while the nodes above represent mathematical operations. GP grows an expression tree in a manner that mimics biological evolution, employing node mutations and subtree crossovers to explore variations in expressions. Candidates are grouped into generations and participate in a tournament selection process, where individuals with the highest fitness scores survive and advance.
Although GP has been successful in discovering human-interpretable solutions for many low-dimensional problems, its discrete combinatorial approach and lengthy search times make it unsuitable for large and high-dimensional datasets.
An alternative approach to SR involves using a DL framework, such as training a neural network (NN) with activation functions that generalize to broader mathematical operations, including unary functions like (($\cdot$( ⋅)2 and sin($\cdot$), as well as binary functions like $+$ and $\times$. The NN is trained with enforced sparse connections, ensuring that the final expressions derived from the NN are compact enough to be human-interpretable or efficiently deployable in resource-constrained environments.
@@
In addition to benefiting from faster gradient-based optimization, the DL approach can utilize GPUs to accelerate both training and inference, whereas GP-based algorithms are typically limited to CPUs. The key to training an NN that produces effective SR results, balancing model performance and complexity, is controlling sparsity. However, most recent developments in using NNs for SR have relied on less efficient pruning methods to achieve the necessary sparsity, requiring multiple training phases with hard-threshold pruning followed by fine-tuning. These multistage frameworks often result in significant performance compromise, as accuracy and sparsity are optimized in separate training phases. The lack of an integrated sparsity control scheme prevents DL approaches from fully realizing their potential to expand SR’s applicability to a broader range of problems.
@@
In this contribution, we introduce $\tt{SymbolNet}$ [^2] A tensorflow [3] implementation code is available at https://github.com/hftsoi/SymbolNet, a DL approach to SR utilizing NN in a novel and SR-dedicated pruning framework, with the following properties:
@@
As far as we are aware, most of the SR literature has primarily tested their methods on datasets with input dimensions below $\mathcal{O}(10)$. These methods have yet to be demonstrated as efficient solutions for high-dimensional problems such as MNIST and beyond. We validate our framework by learning compact expressions from datasets with input dimensions ranging from $\mathcal{O}(10)$ to $\mathcal{O}(1000)$, demonstrating the effectiveness of the model in solving more practical problems for deployment with constrained computational resources.
@@ 2 Related work
Approaches to SR have traditionally been based on GP, first formulated in [5], arising from the idea of creating a program that enables a computer to solve problems in a manner similar to natural selection and genetic evolution. $\tt{Eureqa}$ [6] is one of the first GP-based SR libraries, but it was developed as a commercial proprietary tool, limiting its accessibility for scientific research. $\tt{PySR}$ [7] is a recently developed open-source library built upon the classic GP approach, which improves with a novel evolve-simplify-optimize loop, making it suitable for practical SR and the automatic discovery of scientific equations [8, 9, 10, 11, 12, 13, 14]. Other examples include $\tt{gplearn}$ [15], $\tt{Operon}$ [16], and $\tt{GP-GOMEA}$ [17].
@@
DL has been highly successful in addressing complex problems in fields such as computer vision and natural language processing [18], yet, its application in the domain of SR has not been thoroughly investigated. Equation learner ($\tt{EQL}$) [19, 20, 21] is one of the first NN architectures proposed for performing SR. The approach involves constructing an NN using primitive mathematical operations for neuron activation, training it to achieve a sparse structure through pruning [22, 23, 24], and then unrolling it to obtain the final expression. To prevent the formation of overly complex equations, a three-stage training scheme is employed. In the first stage, a fully-connected NN is trained without regularization, focusing solely on minimizing regression error to allow the parameters to vary freely and establish a solid starting point. In the second stage, $L_{1}$ regularization is imposed to encourage small weights, leading to the emergence of a sparse connection pattern. Finally, all weights below a certain threshold are set to zero and frozen, effectively enforcing a fixed $L_{0}$ norm. The remaining weights are fine-tuned without any regularization. Later research demonstrated that using the $L_{0.5}^{*}$ regularizer [25, 26], which is constructed from a piecewise function and serves as a smooth variant of $L_{0.5}$, can enforce stronger sparsity than $L_{1}$. These methods were primarily tested on simple dynamic system problems with input dimensions of $\mathcal{O}(1)$.
@@
Other variants of NN-based approaches to SR include $\tt{OccamNet}$ [27], which uses an NN to define a probability distribution over a function space, optimized using a two-step method that first samples functions and then updates the weights so that the probability mass is more likely to produce better-fitting functions. Sparsity is maximized by introducing temperature-controlled softmax layers that sample sparse paths through the NN. $\tt{DSR}$ [28] employs an autoregressive recurrent NN to generate expressions sequentially, optimizing them based on reinforcement learning. $\tt{MathONet}$ [29] is a Bayesian learning framework that incorporates sparsity as priors and is applied to solve differential equations. $\tt{N4SR}$ [30] is a multistage learning framework that allows integration of domain-specific prior knowledge. Another class of approaches utilizes transformers pre-trained on large-scale synthetic datasets to generate symbolic expressions from data. For those interested, further details can be found in [31, 32, 33, 34].
@@ 3 SymbolNet architecture
In this section, we describe the model architecture and training framework. We construct an NN composed of symbolic layers, using generic mathematical operators as activation functions. Trainable thresholds are introduced to dynamically prune model weights, input features, and operators. Additionally, a self-adaptive regularization term is introduced for each pruning type to ensure convergence to the target sparsity ratio.
@@ 3.2 Dynamic pruning per network component type
Inspired by this method, we adapt it in our model to perform a weight-wise pruning. We then generalize the idea to also prune input features and mathematical operators, introducing a regularization strategy to achieve convergence to the desired sparsity ratios.
To implement dynamic pruning, Ref. [4] used a step function as a masking function, along with a piecewise polynomial estimator that is nonzero and finite in the range $[-1,1]$ but zero elsewhere. This approach was used to approximate the derivative in the backward pass, preventing a zero gradient in the thresholds, as the original derivative is almost zero everywhere. In this work, we employ a similar binary step function $\theta(x)=\mathbf{1}_{x>0}$ (i.e., 1 if $x>0$, otherwise 0) for masking in the forward pass, along with a smoother estimator in the backward pass using the derivative of the sigmoid function $\frac{d\theta(x)}{dx}\approx\frac{\kappa e^{-\kappa x}}{(1+e^{-\kappa x})^{2}}$ with $\kappa=5$, which is nonzero everywhere.
@@ 3.2.2 Pruning of input features
For the input vector $\bm{x}\in\mathbb{R}^{n_{\text{input}}}$, there are an auxiliary weight vector and a threshold vector $\bm{W}_{\text{input}},\bm{T}_{\text{input}}\in\mathbb{R}^{n_{\text{input}}}$. The pruning operation is implemented using the step function vector $\bm{\Theta}_{\text{input}}=\theta(\bm{W}_{\text{input}}-\bm{T}_{\text{input}})$, applied as follows: $x_{i}\rightarrow(\bm{x}\odot\bm{\Theta_{\text{input}}})_{i}=x_{i}\theta(1-T_{\text{input},i})$, as illustrated in Fig. 3b. In other words, an input node is pruned when its associated threshold value reaches 1.
@@ 3.2.3 Pruning of mathematical operators
Similarly, for each binary operator $g(\cdot,\cdot)$: $\mathbb{R}^{2}\rightarrow\mathbb{R}$, we associate it with an untrainable auxiliary weight fixed at $w_{\text{binary}}=1$ and a trainable pruning threshold $t_{\text{binary}}\in[0,1]$. The pruning operation is applied to each binary operator node as $g(\cdot,\cdot)\rightarrow[g(\cdot,\cdot)\theta(1-t_{\text{binary}})+\text{Add}(\cdot,\cdot)(1-\theta(1-t_{\text{binary}}))]$, where $\text{Add}(\cdot,\cdot)$ is the addition operator, as illustrated in Fig. 3d. This means a binary operator will be simplified to an addition operator when necessary. We denote all thresholds of this type by a vector $\bm{T}_{\text{binary}}$, with a dimension of $n_{\text{binary}}$, representing the total number of binary operators in the architecture.
@@ 3.4 Training framework
In this single-phase training framework, the overall sparse structure is divided into sparse substructures for the model weights, input features, unary operators, and binary operators, respectively. In particular, a set of sparse input features is automatically searched without the need for an external feature selection process. Each of these sparse substructures is dynamically determined by the competition between the corresponding weights and thresholds. Furthermore, the sparse structure dynamically competes with the regression performance, allowing both to be optimized simultaneously. This approach enables the selection of target sparsity ratios for convergence, rather than the ambiguous tuning of some hyperparameters that lack physical meaning. The final symbolic expressions are derived by unrolling the trained network.
@@ 4.1 Expression complexity
4.2 Baseline for comparison: EQL with multistage pruning
We use the $\tt{EQL}$ architecture [19, 20, 21], trained within a three-stage pruning framework [25], as the baseline for comparison purposes. In this baseline framework, model weights are regularized using the smoothed $L_{0.5}$ term, referred to as $L^{*}_{0.5}$, which is controlled by a free parameter $\lambda$:
@@ 4.2 Baseline for comparison: EQL with multistage pruning
where we set $a=0.01$. In the first phase, $L_{0.5}^{*}$ is turned off ($\lambda=0$) to allow the model to fully learn the regression and establish a solid starting point for the model weights. In the second phase, $L_{0.5}^{*}$ is activated ($\lambda>0$) to reduce the magnitudes of the weights, promoting the emergence of a sparse model structure. After this, weights with small magnitudes are set to zero and subsequently frozen. In the final phase, $L_{0.5}^{*}$ is turned off again ($\lambda=0$) during the fine-tuning of the sparse model.
In the experiments that follow, we configure the baseline with network sizes and operator choices similar to those used for $\tt{SymbolNet}$. We vary $\lambda$ from $10^{-4}$ to $10^{-1}$ and adjust the hard pruning threshold from $10^{-4}$ to $10^{-1}$ to conduct a complexity scan by running multiple trials.

### v3

@@ Abstract
Compact symbolic expressions have been shown to be more efficient than neural network models in terms of resource consumption and inference speed when implemented on custom hardware such as FPGAs, while maintaining comparable accuracy [1]. These capabilities are highly valuable in environments with stringent computational resource constraints, such as high-energy physics experiments at the CERN Large Hadron Collider. However, finding compact expressions for high-dimensional datasets remains challenging due to the inherent limitations of genetic programming, the search algorithm of most symbolic regression methods. Contrary to genetic programming, the neural network approach to symbolic regression offers scalability to high-dimensional inputs and leverages gradient methods for faster equation searching. Common ways of constraining expression complexity often involve multistage pruning with fine-tuning, which can result in significant performance loss. In this work, we propose $\tt{SymbolNet}$, a neural network approach to symbolic regression specifically designed as a model compression technique, aimed at enabling low-latency inference for high-dimensional inputs on custom hardware such as FPGAs. This framework allows dynamic pruning of model weights, input features, and mathematical operators in a single training process, where both training loss and expression complexity are optimized simultaneously. We introduce a sparsity regularization term for each pruning type, which can adaptively adjust its strength, leading to convergence at a target sparsity ratio. Unlike most existing symbolic regression methods that struggle with datasets containing more than $\mathcal{O}(10)$ inputs, we demonstrate the effectiveness of our model on the LHC jet tagging task (16 inputs), MNIST (784 inputs), and SVHN (3072 inputs).
@@ 1 Introduction
Symbolic regression (SR) is a supervised learning method that searches for analytic expressions that best fit the data (see Ref.[2] for recent efforts). Unlike traditional regression methods, such as linear and polynomial regression, SR can model a much broader range of complex datasets because it does not require a pre-defined functional form, which itself is dynamically evolving in the fit.
@@
By expressing models in symbolic forms, SR facilitates human interpretation of the data, enabling the potential inference of underlying principles governing the observed system, in contrast to the opaque nature of black-box deep learning (DL) models. A historical example is Max Planck’s 1900 empirical fitting of a formula to the black-body radiation spectrum [3], known as Planck’s law. This symbolically fitted function not only inspired the physical derivation of the law but also played a key role in the revolutionary development of quantum theory. Moreover, due to its compact representation compared to most DL models, SR can also be used as a distillation method for model compression [1]. This can accelerate inference time and reduce computational costs, making it particularly valuable in resource-constrained environments. However, a significant drawback of SR is its inherent complexity. The search space for equations expands exponentially with the number of building blocks (variables, mathematical operators, and constants), making it a challenging combinatorial problem. In fact, finding the optimal candidate has been shown to be NP-hard [4].
@@
Genetic programming (GP) has traditionally been the primary approach to SR [5, 6, 7, 8, 9]. It constructs expressions using a tree representation, where the algebraic relations are reflected in the tree’s structure. The tree’s lowest nodes consist of constants and variables, while the nodes above represent mathematical operations. GP grows an expression tree in a manner that mimics biological evolution, employing node mutations and subtree crossovers to explore variations in expressions. Candidates are grouped into generations and participate in a tournament selection process, where individuals with the highest fitness scores survive and advance. Although GP has been successful in discovering human-interpretable solutions for many low-dimensional problems, its discrete combinatorial approach and lengthy search times make it unsuitable for large and high-dimensional datasets.
@@
An alternative approach to SR involves using a DL framework [10, 11, 12, 13, 14, 15, 16, 17, 18], such as training a neural network (NN) with activation functions that generalize to broader mathematical operations, including unary functions like (($\cdot$( ⋅)2 and sin($\cdot$), as well as binary functions like $+$ and $\times$. The NN is trained with enforced sparse connections, ensuring that the final expressions derived from the NN are compact enough to be human-interpretable or efficiently deployable in resource-constrained environments. In addition to benefiting from faster gradient-based optimization, the DL approach can utilize GPUs to accelerate both training and inference, whereas GP-based algorithms are typically limited to CPUs. The key to training a NN that produces effective SR results, balancing model performance and complexity, is controlling sparsity. However, most recent developments in using NNs for SR have relied on less efficient pruning methods to achieve the necessary sparsity, requiring multiple training phases with hard-threshold pruning followed by fine-tuning. These multistage frameworks often result in significant performance compromise, as accuracy and sparsity are optimized in separate training phases. The lack of an integrated sparsity control scheme prevents DL approaches from fully realizing their potential to expand SR’s applicability to a broader range of problems.
@@
In this contribution, we introduce $\tt{SymbolNet}$ [^2] A tensorflow [19] implementation code is available at https://github.com/hftsoi/SymbolNet, a DL approach to SR utilizing NN in a novel and SR-dedicated pruning framework, specifically designed as a model compression technique for low-latency inference applications, with the following properties:
@@
As far as we are aware, most of the SR literature has primarily tested their methods on datasets with input dimensions below $\mathcal{O}(10)$. These methods have yet to be demonstrated as efficient solutions for high-dimensional problems such as MNIST and beyond. Our framework leverages the compact representation of symbolic expressions and is designed to enable low-latency inference on custom hardware such as FPGAs, targeting high-dimensional datasets without ground-truth equations, as commonly encountered in high-energy experiments at the CERN Large Hadron Collider (LHC). We validate our framework by learning compact expressions from datasets with input dimensions ranging from $\mathcal{O}(10)$ to $\mathcal{O}(1000)$, demonstrating the effectiveness of the model in solving more practical problems for deployment with constrained computational resources.
@@ 2 Related work
SR has started to gain significant attention over the past decade [21, 22, 23, 2]. Approaches to SR have traditionally been based on GP, first formulated in [24], arising from the idea of creating a program that enables a computer to solve problems in a manner similar to natural selection and genetic evolution. $\tt{Eureqa}$ [5] is one of the first GP-based SR libraries, but it was developed as a commercial proprietary tool, limiting its accessibility for scientific research. $\tt{PySR}$ [6] is a recently developed open-source library built upon the classic GP approach, augmented with a novel evolve-simplify-optimize loop, making it suitable for practical SR and the automatic discovery of scientific equations [25, 26, 27, 28, 29, 30, 31]. Other examples include $\tt{gplearn}$ [7], $\tt{Operon}$ [8], and $\tt{GP-GOMEA}$ [9]. For scalable SR, feature selection can be employed prior to performing the equation search. For instance, $\tt{PySR}$ integrates an external random forest regression algorithm to handle high-dimensional datasets by selecting a subset of features based on their relevance to a regression task, which are then passed to the main GP algorithm. However, this approach is practically viable only if the number of selected features remains low, such as below ten, as the search algorithm is still being performed with GP. Alternative methods include the fast function extraction [32, 33] and differentiable GP approach [34], which aim to address the scalability challenges of SR but are not NN-based.
@@
DL has been highly successful in addressing complex problems in fields such as computer vision and natural language processing [35], yet, its application in the domain of SR has not been thoroughly investigated. Equation learner ($\tt{EQL}$) [10, 11, 12] is one of the first NN architectures proposed for performing SR. The approach involves constructing a NN using primitive mathematical operations for neuron activation, training it to achieve a sparse structure through pruning [36, 37, 38], and then unrolling it to obtain the final expression. To prevent the formation of overly complex equations, a three-stage training scheme is employed. In the first stage, a fully-connected NN is trained without regularization, focusing solely on minimizing regression error to allow the parameters to vary freely and establish a solid starting point. In the second stage, $L_{1}$ regularization is imposed to encourage small weights, leading to the emergence of a sparse connection pattern. Finally, all weights below a certain threshold are set to zero and frozen, effectively enforcing a fixed $L_{0}$ norm. The remaining weights are fine-tuned without any regularization. Later research demonstrated that using the $L_{0.5}^{*}$ regularizer [13, 14], which is constructed from a piecewise function and serves as a smooth variant of $L_{0.5}$, can enforce stronger sparsity than $L_{1}$. These methods were primarily tested on simple dynamic system problems with input dimensions of $\mathcal{O}(1)$.
@@
Other variants of NN-based approaches to SR include $\tt{OccamNet}$ [15], which uses a NN to define a probability distribution over a function space, optimized using a two-step method that first samples functions and then updates the weights so that the probability mass is more likely to produce better-fitting functions. Sparsity is maximized by introducing temperature-controlled softmax layers that sample sparse paths through the NN. $\tt{DSR}$ [16] employs an autoregressive recurrent NN to generate expressions sequentially, optimizing them based on reinforcement learning. $\tt{MathONet}$ [17] is a Bayesian learning framework that incorporates sparsity as priors and is applied to solve differential equations. $\tt{N4SR}$ [18] is a multistage learning framework that allows integration of domain-specific prior knowledge. Another class of approaches utilizes transformers pre-trained on large-scale synthetic datasets to generate symbolic expressions from data. For those interested, further details can be found in [39, 40, 41, 42].
@@ 3 SymbolNet architecture
In this section, we describe the model architecture and training framework. We construct a NN composed of symbolic layers, using generic mathematical operators as activation functions. Trainable thresholds are introduced to dynamically prune model weights, input features, and operators. Additionally, a self-adaptive regularization term is introduced for each pruning type to ensure convergence to the target sparsity ratio.
@@ 3.2 Dynamic pruning per network component type
Inspired by this method, we incorporate it into our model to perform a weight-wise pruning. We then generalize the idea to also prune input features and mathematical operators, introducing a regularization strategy to achieve convergence to the desired sparsity ratios.
To implement dynamic pruning, Ref. [20] used a step function as a masking function, along with a piecewise polynomial estimator that is nonzero and finite in the range $[-1,1]$ but zero elsewhere. This approach was used to approximate the derivative in the backward pass, preventing a zero gradient in the thresholds, since the derivative of the original step function is almost zero everywhere. In this work, we employ a similar binary step function $\theta(x)=\mathbf{1}_{x>0}$ (i.e., 1 if $x>0$, otherwise 0) for masking in the forward pass, along with a smoother estimator in the backward pass using the derivative of the sigmoid function $\frac{d\theta(x)}{dx}\approx\frac{\kappa e^{-\kappa x}}{(1+e^{-\kappa x})^{2}}$ with $\kappa=5$ (a very high $\kappa$ would effectively turn it into a delta function, while a very low $\kappa$ would make it flat and less sensitive to the step location), which is nonzero everywhere.
@@ 3.2.2 Pruning of input features
For the input vector $\bm{x}\in\mathbb{R}^{n_{\text{input}}}$, there is an auxiliary weight vector and a threshold vector $\bm{W}_{\text{input}},\bm{T}_{\text{input}}\in\mathbb{R}^{n_{\text{input}}}$. The pruning operation is implemented using the step function vector $\bm{\Theta}_{\text{input}}=\theta(\bm{W}_{\text{input}}-\bm{T}_{\text{input}})$, applied as follows: $x_{i}\rightarrow(\bm{x}\odot\bm{\Theta_{\text{input}}})_{i}=x_{i}\theta(1-T_{\text{input},i})$, as illustrated in Fig. 3b. In other words, an input node is pruned when its associated threshold value reaches 1.
@@ 3.2.3 Pruning of mathematical operators
Similarly, for each binary operator $g(\cdot,\cdot)$: $\mathbb{R}^{2}\rightarrow\mathbb{R}$, we associate it with an untrainable auxiliary weight fixed at $w_{\text{binary}}=1$ and a trainable pruning threshold $t_{\text{binary}}\in[0,1]$. The pruning operation is applied to each binary operator node as $g(\cdot,\cdot)\rightarrow[g(\cdot,\cdot)\theta(1-t_{\text{binary}})+\text{Add}(\cdot,\cdot)(1-\theta(1-t_{\text{binary}}))]$, where $\text{Add}(\cdot,\cdot)$ is the addition operator, as illustrated in Fig. 3d. This means a binary operator will be simplified to an addition operator when necessary, which is motivated by the fact that addition is typically simpler to compute than other more complex binary operators. For instance, an addition operation requires only one clock cycle to execute on an FPGA, while other binary operators may generally take significantly longer. We denote all thresholds of this type by a vector $\bm{T}_{\text{binary}}$, with a dimension of $n_{\text{binary}}$, representing the total number of binary operators in the architecture.
@@ 3.4 Training framework
In this single-phase training framework, the overall sparse structure is divided into sparse substructures for the model weights, input features, unary operators, and binary operators, respectively. In particular, a set of sparse input features is automatically searched without the need for an external feature selection process. Each of these sparse substructures is dynamically determined by the competition between the corresponding weights and thresholds. Furthermore, the sparse structure dynamically competes with the regression performance, allowing both to be optimized simultaneously. This approach enables direct specification of target sparsity ratios for each component (weights, features, and operators), rather than requiring the indirect tuning of multiple regularization parameters whose relationships to the resulting sparsity levels are more difficult to predict. The final symbolic expressions are derived by unrolling the trained network.
@@ 4.1 Expression complexity
4.2 Baseline for comparison
From DL-based methods, we use the $\tt{EQL}$ architecture [10, 11, 12], trained within a three-stage pruning framework [13], as the baseline for comparison purposes. In this baseline framework, model weights are regularized using the smoothed $L_{0.5}$ term, referred to as $L^{*}_{0.5}$, which is controlled by a free parameter $\lambda$:
@@ 4.2 Baseline for comparison
where we set $a=0.01$. In the first phase, $L_{0.5}^{*}$ is turned off ($\lambda=0$) to allow the model to fully learn the regression and establish a solid starting point for the model weights. In the second phase, $L_{0.5}^{*}$ is activated ($\lambda>0$) to reduce the magnitudes of the weights, promoting the emergence of a sparse model structure. After this, weights with small magnitudes are set to zero and subsequently frozen. In the final phase, $L_{0.5}^{*}$ is turned off again ($\lambda=0$) during the fine-tuning of the sparse model. In the experiments that follow, we configure the baseline with network sizes and operator choices similar to those used for $\tt{SymbolNet}$. We vary $\lambda$ from $10^{-4}$ to $10^{-1}$ and adjust the hard pruning threshold from $10^{-4}$ to $10^{-1}$ to conduct a complexity scan by running multiple trials.
Additionally, we use the GP-based $\tt{PySR}$ as another baseline for comparison. In general, the efficiency and performance of GP-based SR algorithms tend to degrade as the number of input features increases, typically beyond ten. To address this, $\tt{PySR}$ incorporates an external feature selector based on a gradient-boosting tree algorithm to identify important features before feeding into the main search loop, particularly for datasets with high input dimensionality [6]. For the LHC dataset, which has 16 input features, we configures $\tt{PySR}$ to select only half of features before performing equation searches. However, for MNIST (784 features) and SVHN (3072 features), which feature counts exceeding the LHC dataset by more than an order of magnitude, reducing the number of features to fewer than 10 makes the models less accurate, while exceeding 10 makes GP-based searches computationally impractical, especially since the outputs are multidimensional as well. Therefore, we restrict our comparison with $\tt{PySR}$ to the LHC dataset only.